# Statistical Principles for Platform Trials

### v4

@@ 3.1.1 Effect of correlation induced by sharing control on variability of $V^{*}$
In platform trials, there are several methods to compute the covariances of non-coverage events across different treatment arms. Specifically, let $T^{\star}_{i}$ represent the pivoted statistic used to compare the new compound $i$, $i\in N$, against a shared control in a platform trial, such that
$T^{\star}_{i}=\dfrac{\hat{\mu}_{i}-\hat{\mu}_{SC}}{SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}}$ (3.1)
@@
has a null distribution free of the efficacy parameter, where $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ are estimated mean of the new compound $i$ and the shared control, respectively. For proof of concept, consider the case where each pair $(T^{\star}_{i},T^{\star}_{j})$, with $i\neq j$, follows a bivariate standard Normal distribution. This scenario is particularly applicable in therapeutic areas like Alzheimer’s Disease (AD) and Type 2 Diabetes Mellitus (T2DM), where the primary efficacy endpoints11footnotetext: which excludes the score test version of the logRank test are well-defined and standardized. The covariance between non-coverage indicators $Z_{i}$ and $Z_{j}$ is given by:
$Cov(Z_{i},Z_{j})=P\{T^{\star}_{i}>z_{\alpha}\cap T^{\star}_{j}>z_{\alpha}\}-\alpha^{2},$ (3.2)
which depends on $\rho_{ij}=Corr(T^{\star}_{i},T^{\star}_{j})=\lambda_{i}\lambda_{j}$, where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$.
@@
Each subset $N$ has its distinct $\rho_{ij}$ matrix, and thus $\lambda$ matrices, depending not only on the ratio of sample sizes between the treatment arms with index in $N$ and the shared control arm, but also on other covariates like baseline measurement of each patient. To assess the stability of error rate control for compounds apparently with no efficacy, we consider the observed treatment outcomes as well as the covariates as known. Consequently, we compute $\rho_{ij}=\lambda_{i}\lambda_{j},i\neq j,$ among the compounds in $N$ that show no efficacy. Such computations are feasible within model-based analytical frameworks.
Table 1 illustrates that the standard deviation of $V^{\star}$, denoted as $StdDev(V^{\star})$, increases with higher correlation coefficients $\rho_{ij}$. If the treatment arms were independent, $StdDev(V^{\star})$ exceeds the mean $E(V^{\star})$. Moreover, correlations among the test statistics, caused by sharing a control, further increase the variability of $V^{\star}$ (the number of incorrect approvals). For example, a common correlation of 0.5 among the test statistics raises $StdDev(V^{\star})$ from 0.69 to 1.16. This inherent variability of the actual error rate in a specific family of clinical trials is a significant issue that, we believe, warrants greater attention in practice, irrespective of whether the error rate is measured as ERpF or ERFw.
$k$ $E(V^{\star}$ ) $StdDev(V^{\star})$
Unconditional Independent studies Hypothetical Platform $k:1$ randomization
$n^{\mathrm{true}}$ ERpF $\rho_{ij}=0$ $\rho_{ij}=0.3$ $\rho_{ij}=0.5$
5 0.25 0.49 0.57 0.66
10 0.5 0.69 0.94 1.16
20 1 0.97 1.65 2.15
40 2 1.38 3.02 4.12
Table 1: For a family of studies, if $StdDev(V^{\star})>E(V^{\star})$, that is, standard deviation of the number of errors is larger than the mean of the number of errors, then one might consider stability of the ERpF control to be of concern. In a Platform trial, pairs of $T^{\star}_{i}$ and $T^{\star}_{j}$ are correlated due to sharing of Control patients. Assuming $\rho_{ij}$ is common for all pairs, variability of $V^{\star}$ increases with increasing $\rho_{ij}$. For illustration, consider a Platform trial with $k$ new compounds using a $k:1$ compound-to-placebo randomization (which makes the sample size for comparing each new compound with the shared control $1:1$). With ERpF set at $5\%\times n^{\mathrm{true}}$ where $n^{\mathrm{true}}$ is the number of in-efficacious compounds, we see independent studies reach stability by family size 20, but that is not the case for studies in platform trials. Stability of ERpF control is thus more of a concern in platform trials.
To facilitate the assessment of the impact of correlations, analogous to how internet commerce tracks cookies, we propose maintaining an informational “cookie” for each treatment or regimen in a platform trial.
Control Sharing Matrix
Each treatment or regimen should have a “cookie” that records the exact time interval of its enrollment period. By identifying the overlap between the enrollment periods of different regimens, it becomes possible to compute the pairwise covariance and correlation for comparisons between the active treatment and the control.
$n$ A B C D
A 122 122 164 103
B 122 120 122 103
C 164 122 120 103
D 103 103 103 75
Table 2: Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial
A B C D
A 1 0.498 0.425 0.478
B 0.498 1 0.496 0.476
C 0.425 0.496 1 0.476
D 0.478 0.476 0.476 1
Table 3: Correlations matrix between $\hat{\mu}_{i}-\hat{\mu}_{SC_{ij}}$ and $\hat{\mu}_{j}-\hat{\mu}_{SC_{ij}}$
for the HEALEY ALS Platform Trial
A B C D
A 0.0475 0.0093 0.0083 0.0087
B 0.0093 0.0475 0.0092 0.0096
C 0.0083 0.0092 0.0475 0.0086
D 0.0087 0.0096 0.0086 0.0475
Table 4: Matrix of variances of $Z_{i}$ (diagonal) and covariances between $Z_{i}$ and $Z_{j}$ (off-diagonal) with ERpF set at $5\%\times 4=0.2$ for Regimens ABCD of the ALS Platform trial.
@@ A B C D
Table 2 presents the Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial, inferred from the results available on clinicaltrials.gov. For each regimen, the diagonal elements represent the number of patients randomized to the active treatment. The off-diagonal elements indicate the number of patients shared in the control group, to which each active treatment is compared, for each pair of regimens. We can then derive Table 3, which presents the correlation matrix (Details of the derivation are available in the supplementary materials). Note that as mentioned before, the pairwise correlation in Table 3 can also be expressed as $\hat{\rho}_{ij}=\lambda_{i}\lambda_{j}$ for all $i<j\in N$, where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$. These $\lambda$ values can be readily calculated using existing software, making this notation preferable for its ease of extension to more complex structures. For the HEALEY ALS Platform Trial, the approximated $\lambda=$ (0.674, 0.722, 0.671, 0.690). Based on equation (3.2), we can obtain variance-covariance matrix for $Z$ (see Table 4). Note that the elements sum to 0.2974 for $Var(V^{\star})$, resulting in $StdDev(V^{\star})$ = 0.5454, which exceeds the ERpF of 0.2.
3.1.2 Effect of sample mean of shared control on variability of $V^{*}$
Assuming, without loss of generality, that $\mu_{i}=0$ for $i\in N$. According to equation (3.1), a positive (or negative) estimated mean $\hat{\mu}_{SC}$ of the shared control can make it harder (or easier) for new compounds to obtain approval. In this paper, a positive $\hat{\mu}_{SC}$ is referred to as an under-performing estimated mean of the shared control, while a negative $\hat{\mu}_{SC}$ is termed an over-achieving estimated mean of the shared control.
@@ 3.1.2 Effect of sample mean of shared control on variability of $V^{*}$
An under-performing $\hat{\mu}_{SC}$ presents significant concerns for sponsors seeking treatment approvals. The higher $\hat{\mu}_{SC}$ becomes, the more challenging it is for treatment arms to receive approval. This scenario leads to high probability of $V^{\star}$ being towards lower numbers of erroneous approvals, resulting in a lower variability of $V^{\star}$. However, for regulators, especially when the treatment arms are non-efficacious, this situation is less problematic as it reduces the likelihood of erroneously approving a treatment.
@@
With over-achieving $\hat{\mu}_{SC}$, the situation becomes a concern for regulators, as it is much easier to approve treatment arms due to the lower $\hat{\mu}_{SC}$. This increased ease of approval heightens the risk of erroneous approvals. In extreme cases, if $\hat{\mu}_{SC}$ is sufficiently low, almost all treatment arms may be erroneously approved. While sponsors might favor such an outcome, it is undesirable from a patients and regulatory perspective because it leads to the approval of non-efficacious drugs, and increases the probability of higher $V^{\star}$.
These two cases are demonstrated in Figure (1). Assume $\mu_{i}=0$ for $i=1,2,3,4\in N$, and $\alpha=0.05$. Then, Figure (1) compares the true probability distribution of $V^{*}$ when $\hat{\mu}_{SC}={\mu}_{SC}=0$ (the dash line) with those where $\hat{\mu}_{SC}$ is either greater than or less than ${\mu}_{SC}=0$ (the solid line). Figure 1a indicates an essentially negligible change in the probability of obtaining $V^{*}=1$, and the probabilities of obtaining $V^{*}=2,3,4$ are nearly zero, suggesting minimal impact on the variability of $V^{*}$. However, as shown in Figures 1b and 1c, the likelihood of obtaining $V^{*}=1$ and $V^{*}=2$ increases as $\hat{\mu}_{SC}$ decreases, thereby increasing the variability of $V^{*}$.
(a) $\hat{\mu}_{SC}=0.46$
1 SE above $\mu_{SC}$
(b) $\hat{\mu}_{SC}=-0.23$
0.5 SE below $\mu_{SC}$
(c) $\hat{\mu}_{SC}=-0.46$
1 SE below $\mu_{SC}$
Figure 1: Probability distribution of $V^{\star}$. In this scenario, we have 5 treatment arms, each with an $\alpha$ level of 0.05, resulting in an ERpF of $0.05\times 5=0.25$. The total sample size is 1200, with a 5:1 ratio of treatment arm to control sample size allocation, leading to an unconditional correlation of 0.5 between estimates in differences between each new treatment arm and the control. For illustration purpose, we assume that the true shared control mean $\mu_{SC}$ and treatment arm means $\mu_{i}$ both equal to 0 (although these values can be adjusted). In Figure 1(a), the distribution of $V^{\star}$ closely mirrors the underlying true distribution. With a higher $\hat{\mu}_{SC}$, it becomes more difficult for treatment arms to gain approval, leading to a negligible probability of $V^{\star}>0$. Conversely, Figure 1(b) and 1(c) show scenarios where the $\hat{\mu}_{SC}$ is much lower than $\mu_{i}$, facilitating easier erroneous approvals. In particular, a substantial increase for $V^{\star}=1$ and $V^{\star}=2$ would not cause only an increase in $V^{\star}$, but also its variability.
3.1.3 Effect of mean and correlation induced by sharing control on variability of $V^{*}$
It is to be noted that while both the mean and the correlation induced by sharing control individually affect the distribution of $V^{\star}$, it is important to consider their combined effect on the distribution of $V^{\star}$. Specifically, the variance of $V^{\star}$ is noteworthy. We standardized $\hat{\mu}_{SC}$ by $SE_{SC}$ which we denote by $\hat{\mu}_{SC}^{\star}$, i.e.
$\hat{\mu}_{SC}^{\star}=\frac{\hat{\mu}_{SC}}{SE_{SC}}.$
As seen in Figure 2, for cases when $\hat{\mu}_{SC}^{\star}$ is greater than the treatment arm mean $\hat{\mu}_{i}$, the performance of the standard deviation of $V$ is lower and much more stable, which translates to a more stable $V^{\star}$ value. On the other hand, when $\hat{\mu}_{SC}^{\star}$ is less than $\hat{\mu}_{i}$, the behavior of the standard deviation of $V$ is higher and much more variable, which translates to a more variable $V^{\star}$. Due to this behavior of $V^{\star}$, it is beneficial to look at the whole behavior of the probability distribution of $V$ rather than single instances of $V$.
Figure 2: The behavior of standard deviation for $V^{\star}$ is heavily affected by correlation and $\hat{\mu}_{SC}$. Assume $\mu_{SC}=0$, i.e. it is non efficacious. As we decrease the value of our $\hat{\mu}_{SC}$ and thus $\hat{\mu}_{SC}^{\star}$, the performance of standard deviation of $V^{\star}$ becomes less stable.
$\hat{\mu}_{SC}^{\star}$ = -3
$\hat{\mu}_{SC}^{\star}$ = -1.5
$\hat{\mu}_{SC}^{\star}$ = 0
$\hat{\mu}_{SC}^{\star}$ = 1.5
$\hat{\mu}_{SC}^{\star}$ = 3
Figure 3: Snapshot of different standardized $\hat{\mu}_{SC}$ values on correlation’s effect on standard deviation where $\mu_{i}=0$. As can be seen, when the $\hat{\mu}_{SC}^{\star}$ is less than the treatment mean, the behavior is less simple whereas for $\hat{\mu}_{SC}^{\star}$ is greater than the treatment mean, the standard deviation of V decreases as correlation increases.
3.2 Conditional assessment of FAR control
A foundational principle in (frequentist) Statistics is
Principle 3.1 (The Conditionality Principle).
See J. O. Berger (1985) page 31 and Casella and R. L. Berger (1990) page 268 for details.
To motivate the application of this principle to conditional confidence, consider Fieller’s confidence interval for the ratio of the means of two Normal distributions (Fieller 1954) which unconditionally (over infinitely many replications of a study) has confidence level exactly equal to $(1-\alpha)$. Interestingly, depending on the data observed in each particular study, Fieller’s confidence interval can be an interval, or the entire real line, the complement of an interval, or the empty set (see Casella and Berger 1990, page 459). To us, conditional on Fieller’s interval being the entire real line for a particular study, a conditional confidence of 100% should be stated for that study. Likewise, conditional on Fieller’s interval being the empty set for a particular study, a conditional confidence of 0% should be stated for that study, not averaging the coverage probability over studies that have not been performed that would result in Fieller’s confidence interval being an interval, the complement of one, or the entire real line. In this spirit, while taking comfort that a 5% ERFw within each study results in an unconditional ERpF of 5% across all the studies, we recommend assessing the conditional ERpF within each Platform trial conditional on the observed shared Control outcome, to be sure they do not deviate too much from the unconditional ERpF.
4 A Confident Platform Random Sizer App

### v5

@@ 3.3.1 Effect of correlation induced by sharing control on variability of $V^{*}$
Suppose each confidence interval is constructed from a pivotal statistic
$T^{\star}_{i}=\dfrac{\hat{\mu}_{i}-\hat{\mu}_{SC}-\theta_{i}}{SE_{\hat{\mu}_{i}-\hat{\mu}_{SC}}}$ (3.1)
whose sampling distribution does not depend on the efficacy parameter $\theta_{i}$. Here $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ denote the sample mean for new compound arm $i$ and the shared control arm, respectively. For proof of concept, consider the setting where
$\hat{\mu}_{i}\sim N(\mu_{i},\frac{\sigma_{i}^{2}}{n_{i}}),\qquad\hat{\mu}_{SC}\sim N(\mu_{SC},\frac{\sigma_{SC}^{2}}{n_{SC}}),$
Assume further that $\hat{\mu}_{i}$ and $\hat{\mu}_{SC}$ are independent. Then the estimator of the treatment effect is
$\hat{\theta}_{i}=\hat{\mu}_{i}-\hat{\mu}_{SC},$
with variance
$Var(\hat{\theta}_{i})=\frac{\sigma_{i}^{2}}{n_{i}}+\frac{\sigma_{SC}^{2}}{n_{SC}}.$
Consequently, $T^{\star}_{i},i=1,\cdots,k$ have marginal standard normal distribution, and for each pair $i\neq j$,
$(T_{i}^{\star},T_{j}^{\star})\sim N_{2}\!\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho_{ij}\\ \rho_{ij}&1\end{pmatrix}\right)$.
@@
This scenario is particularly applicable in therapeutic areas like Alzheimer’s Disease (AD) and Type 2 Diabetes Mellitus (T2DM), where the primary efficacy endpointsfootnotetext: which excludes the score test version of the logRank test are well-defined and standardized.
Let
$Z_{i}^{\star}=\mathbf{1}\{T_{i}^{\star}>z_{\alpha}\}$
denote the indicator of non-coverage for the one-sided $(1-\alpha)$ confidence interval, where $z_{\alpha}$ is the upper $(1-\alpha)$ quantile of the standard normal distribution. Under this construction,
$P(Z_{i}^{\star}=1)=\alpha.$
Define
$V^{\star}=\sum_{i=1}^{k}Z_{i}^{\star}$
as the total number of non-coverage events across new compound arms. For any $i\in L$, we have $\theta_{i}\leq 0$, and whenever the confidence interval excludes 0 in the efficacy direction (i.e., $Z_{i}=1$), it necessarily does not cover the true value (i.e., $Z_{i}^{*}=1$). Hence,
$Z_{i}\leq Z_{i}^{*}\quad\text{for all }i\in L,$
which implies
$V^{k}=\sum_{i\in L}Z_{i}\;\leq\;\sum_{i\in L}Z_{i}^{*}\;\leq\;\sum_{i=1}^{k}Z_{i}^{*}\;=\;V^{*}.$
Therefore, instead of $V^{k}$, we consider the distribution of $V^{*}$, an upper bound on the number of compounds that were erroneously inferred to be efficacious.
Note that $V^{*}$ counts all one-sided confidence interval noncoverage events in the following cases. Therefore $V^{*}$ is more conservative as a bound.
Case 1: Misses for efficacious arms. Suppose $i\notin L$, so $\theta_{i}>0$. If the confidence interval fails to cover the true value $\theta_{i}$, then
$Z_{i}^{*}=1.$
However, this event does not correspond to a false approval. Thus, it contributes to $V^{*}$ but not to $V^{k}$.
Case 2: False approvals. If $i\in L$ and the confidence interval lies entirely above zero, then
$Z_{i}^{*}=1\quad\text{and}\quad Z_{i}=1,$
which contributes to both $V^{*}$ and $V^{k}$.
The variance of $\operatorname{Var}(V^{\star})$ can be written as
$\operatorname{Var}(V^{\star})=\sum_{i=1}^{k}\operatorname{Var}(Z_{i}^{\star})+2\sum_{1\leq i<j\leq k}\operatorname{Cov}(Z_{i}^{\star},Z_{j}^{\star}),$
where
$\operatorname{Var}(Z_{i}^{\star})=\alpha(1-\alpha).$
Thus the variance of $V^{\star}$ depends on the covariances $\operatorname{Cov}(Z_{i}^{\star},Z_{j}^{\star})$ for $i\neq j$, and we have
$\operatorname{Cov}(Z_{i}^{\star},Z_{j}^{\star})=P(T_{i}^{\star}>z_{\alpha},\;T_{j}^{\star}>z_{\alpha})-\alpha^{2}.$ (3.2)
In platform trials with a shared control arm, the correlation between the standardized treatment effects arises from the common control mean. Assuming independent new compound treatment arms and equal outcome variances, we can easily prove that
$\rho_{ij}=\operatorname{Corr}(T_{i}^{\star},T_{j}^{\star})=\lambda_{i}\lambda_{j},$
where
$\lambda_{i}=\left(1+\frac{n_{SC}}{n_{i}}\right)^{-1/2},$
and $n_{i}$ and $n_{SC}$ denote the sample sizes of new compound treatment arm $i$ and the shared control arm, respectively.
@@
Each subset $L$ has its distinct $\rho_{ij}$ matrix, and thus $\lambda$ matrices, depending not only on the ratio of sample sizes between the treatment arms with index in $L$ and the shared control arm, but also on other covariates like baseline measurement of each patient. To assess the stability of error rate control for compounds apparently with no efficacy, we consider the observed new compound treatment outcomes as well as the covariates as known. Consequently, we compute $\rho_{ij}=\lambda_{i}\lambda_{j},i\neq j,$ among the compounds in $L$ that show no efficacy. Such computations are feasible within model-based analytical frameworks.
Table 1 illustrates how correlation among treatment comparisons affects the variability of $V^{\star}$. As the common correlation $\rho_{ij}$ increases, $StdDev(V^{\star})$ increases substantially and can exceed the mean $E(V^{\star})$. For example, when $k=10$ and $\rho_{ij}=0.5$, $StdDev(V^{\star})$ increases to 1.16 compared with 0.69 under independence. Moreover, the variability grows rapidly with family size $k$. If $StdDev(V^{\star})>E(V^{\star})$, the realized number of erroneous claims within a given family may fluctuate widely around its expectation, raising concerns about the stability of ERpF control. While independent studies approach stability when the family size is around $k=20$, this stabilization does not occur in platform trials due to the correlation induced by the shared control arm.
@@
To facilitate the assessment of the impact of correlations, analogous to how internet commerce tracks cookies, we propose maintaining an informational “cookie” for each treatment or regimen in a platform trial. Table 2 presents the Control Sharing Matrix (CSM) for the HEALEY ALS Platform Trial, inferred from the results available on clinicaltrials.gov. For each regimen, the diagonal elements represent the number of patients randomized to the active treatment. The off-diagonal elements indicate the number of patients shared in the control group, to which each active treatment is compared, for each pair of regimens. We can then derive Table 3, which presents the correlation matrix. Note that as mentioned before, the pairwise correlation in Table 3 where $\lambda_{i}=(1+\frac{n_{SC}}{n_{i}})^{-1/2}$. These $\lambda$ values can be readily calculated using existing software, making it preferable for its ease of extension to more complex correlation structuresaaaFor example, when two additional regimens are added later on, we do not need to calculate a six by six correlation matrix, instead, we only need to calculate the pairwise correlation based on the corresponding shared control sample size for each pair of regimens. For the HEALEY ALS Platform Trial, the approximated $\lambda=$ (0.674, 0.722, 0.671, 0.690). Based on equation (3.2)bbbEquation (3.2) assumes all the regimens have the same $n_{SC}$, which may not be the case in real life (e.g., the HEALEY ALS Platform Trial). In practice, one just uses the best available approximation, and we use the $\lambda$ from the factor-analytic approximation. we can obtain variance-covariance matrix for $Z^{*}$ (see Table 4). Note that the elements sum to 0.2974 for $Var(V^{\star})$, resulting in $StdDev(V^{\star})$ = 0.5454, which exceeds the ERpF of 0.2.
3.3.2 Effect of sample mean of shared control on the distribution of conditional $V^{*}$
From the perspective of the Conditionality Principle, inference should be evaluated conditional on the realized outcome of shared control, since regulatory decisions are made after this quantity is observed. Recall the standardized statistic
$T_{i}^{\star}=\frac{\hat{\mu}_{i}-\hat{\mu}_{SC}-\theta_{i}}{\sqrt{\frac{\sigma_{i}^{2}}{n_{i}}+\frac{\sigma_{SC}^{2}}{n_{SC}}}},$
for treatment arm $i$, where
$\hat{\mu}_{i}\sim N\!\left(\mu_{i},\frac{\sigma_{i}^{2}}{n_{i}}\right),\quad\hat{\mu}_{SC}\sim N\!\left(\mu_{SC},\frac{\sigma_{SC}^{2}}{n_{SC}}\right),$
and
$\theta_{i}=\mu_{i}-\mu_{SC}.$
Conditional on $\hat{\mu}_{SC}$,
$\hat{\mu}_{i}-\hat{\mu}_{SC}\sim N\!\left(\mu_{i}-\hat{\mu}_{SC},\frac{\sigma_{i}^{2}}{n_{i}}\right),$
and
$\hat{\mu}_{i}-\hat{\mu}_{SC}-\theta_{i}=\hat{\mu}_{i}-\hat{\mu}_{SC}-(\mu_{i}-\mu_{SC}),$
we have
$E(\hat{\mu}_{i}-\hat{\mu}_{SC}-\theta_{i}\mid\hat{\mu}_{SC})=\mu_{SC}-\hat{\mu}_{SC}.$
Let
$m_{i}(\hat{\mu}_{SC})=\frac{\mu_{SC}-\hat{\mu}_{SC}}{\sqrt{\frac{\sigma_{i}^{2}}{n_{i}}+\frac{\sigma_{SC}^{2}}{n_{SC}}}},\qquad v_{i}=\frac{\sigma_{i}^{2}/n_{i}}{\frac{\sigma_{i}^{2}}{n_{i}}+\frac{\sigma_{SC}^{2}}{n_{SC}}},$
we have
$T_{i}^{\star}\mid\hat{\mu}_{SC}\sim N\!\left(m_{i}(\hat{\mu}_{SC}),v_{i}\right).$
Notably, the conditional mean of $T_{i}^{\star}$ does not depend on the efficacy parameter $\theta_{i}$. A larger (or smaller) realized value of $\hat{\mu}_{SC}$ relative to its expectation $\mu_{SC}$ makes it harder (or easier) for experimental compounds to demonstrate efficacy, regardless of the true treatment effect. In this paper, a shared control with $\hat{\mu}_{SC}>\mu_{SC}$ is referred to as an over-achieving shared control, while $\hat{\mu}_{SC}<\mu_{SC}$ is termed an under-performing shared control.
Conditional on the observed shared control mean $\hat{\mu}_{SC}$, the conditional non-coverage probability for arm $i$ is
$p_{i}(\hat{\mu}_{SC}):=P(Z_{i}^{\star}=1\mid\hat{\mu}_{SC})=P(T_{i}^{\star}>z_{\alpha}\mid\hat{\mu}_{SC}).$
so that
$p_{i}(\hat{\mu}_{SC})=1-\Phi\!\left(\frac{z_{\alpha}-m_{i}(\hat{\mu}_{SC})}{\sqrt{v_{i}}}\right),$
where $\Phi(\cdot)$ denotes the cumulative distribution function of the standard normal distribution.
Conditional on $\hat{\mu}_{SC}$, the treatment-arm estimators are independent, and hence $Z_{1}^{\star},\ldots,Z_{k}^{\star}$ are conditionally independent Bernoulli random variables with success probabilities $p_{1}(\hat{\mu}_{SC}),\ldots,p_{k}(\hat{\mu}_{SC})$. Therefore,
$V^{\star}\mid\hat{\mu}_{SC}$
follows a Poisson-binomial distribution in general. If, in addition,
$\sigma_{i}=\sigma_{SC}=\sigma,\qquad n_{i}=n_{SC}=n,\qquad i=1,2,\ldots,k,$
then
$p_{1}(\hat{\mu}_{SC})=\cdots=p_{k}(\hat{\mu}_{SC})=:p(\hat{\mu}_{SC}),$
and therefore
$V^{\star}\mid\hat{\mu}_{SC}\sim\mathrm{Binomial}\bigl(k,p(\hat{\mu}_{SC})\bigr),$
where
$p(\hat{\mu}_{SC})=1-\Phi\!\left(z_{\alpha}\sqrt{2}-\frac{\mu_{SC}-\hat{\mu}_{SC}}{\sigma}\sqrt{n}\right).$
@@ 3.3.2 Effect of sample mean of shared control on the distribution of conditional $V^{*}$
Conditional on an over-achieving $\hat{\mu}_{SC}$, the probability of $V^{\star}|\hat{\mu}_{SC}$ being towards lower numbers of erroneous approvals increases with $\hat{\mu}_{SC}$. In other words, the higher $\hat{\mu}_{SC}$ becomes, the more challenging it is for treatment arms to receive approval, presenting significant concerns for sponsors seeking treatment approvals. However, for regulators, especially when the treatment arms are non-efficacious, this situation is less problematic as it reduces the likelihood of erroneously approving a treatment.
@@
Conditional on an under-performing $\hat{\mu}_{SC}$, the probability of $V^{\star}|\hat{\mu}_{SC}$ being towards high numbers of erroneous approvals increases with decreasing $\hat{\mu}_{SC}$. In extreme cases, if $\hat{\mu}_{SC}$ is sufficiently low, almost all treatment arms may be erroneously approved. While sponsors might favor such increased ease of approval, it is undesirable from a patients and regulatory perspective because it heightens the risk of erroneous approvals.
For illustration, consider a platform trial with $k=5$ treatment arms, each tested at a one-sided significance level $\alpha=0.05$, yielding $\text{ERpF}=0.05\times 5=0.25$. The total sample size is 1200 with a 5:15{:}1 allocation ratio of treatment to control, which induces an unconditional correlation of approximately $0.5$ between the estimated treatment–control differences. For simplicity, we assume the true shared control mean $\mu_{SC}=0$ and the treatment arm means $\mu_{i}=0$ for $i=1,2,3,4,5$. Figure 1 illustrates the conditional distribution of $V^{\star}$ for several realized values of $\hat{\mu}_{SC}$. Figure 1a indicates an essentially negligible change in the probability distribution of $V^{*}|\hat{\mu}_{SC}(\text{over-achieving})$. However, as shown in Figures 1b and 1c, the likelihood of $V^{*}=1|\hat{\mu}_{SC}(\text{under-performing})$ and $V^{*}=2|\hat{\mu}_{SC}(\text{under-performing})$ increases as $\hat{\mu}_{SC}(\text{under-performing})$ decreases, thereby changing the distribution of $V^{*}|\hat{\mu}_{SC}$. These results highlight the importance of examining the full conditional distribution of $V^{*}|\hat{\mu}_{SC}$, rather than relying solely on unconditional error rate or summary statistics.
Figure 1: Probability distribution of $V^{\star}|\hat{\mu}_{SC}$ under a plotform trial with 5 treatment arms, each tested at $\alpha=0.05$, yielding an ERpF of 0.25. The total sample size is 1200, with a 5:1 treatment to control allocation, resulting in an unconditional correlation of 0.5 among treatment-control contrasts. For illustration purpose, the true shared control mean and treatment means are assumed to be $\mu_{SC}=\mu_{i}=0$. Panels (a)-(c) display the conditional distribution of $V^{*}$ given different values of $\hat{\mu}_{SC}$.
3.3.3 A Confident Platform Random Sizer (Interactive visualization via Shiny App)
To facilitate exploration of this conditional distribution, we developed an interactive visualization toolcccCan be found at https://eouya001.shinyapps.io/ConfidentPlatformRandomSizer/ that implements the Conditionality Principle by displaying the entire distribution of $V^{\star}|\hat{\mu}_{SC}$, in line with the FDA draft guidance. Our interactive plot allows users to examine how changes in design parameters affect the entire distribution, highlighting the impact of both over-achieving and under-performing $\hat{\mu}_{SC}$. Users can adjust parameter values interactively and observe in real time how the probabilities of erroneous approvals vary across different realization of $\hat{\mu}_{SC}$, thereby providing a practical way to explore the implications of conditional inference in platform trials.